OpenStudy (anonymous):

How do I expand a logarithm? log4 (3xyz)^2

OpenStudy (anonymous):

Is it 9(log4 x + log4 y + log4 z)?

OpenStudy (anonymous):

log4 (3xyz)^2 = 2 * (log4 (3xyz)) = 2 * (log4 3 + log4 x + log4 y + log4 z) = 2log4 3 + 2log4 x + 2log4 y + 2log4 z)